OpenStudy (anonymous):

whats the inverse of 1/7x-8

OpenStudy (anonymous):

is there anymore to the original statement such as what it =?

OpenStudy (ddcamp):

Let's say y = 1/7x -8 swap the x and y, then solve for y: x = 1/7y - 8 x + 8 = 1/7y 7y = 1/(x+8) y = 1/7(x+8)